OpenStudy (anonymous):

What are the zeros of the polynomial function: f(x) = x3 + 4x2 + x – 6 x = –3, x = –1, x = 2 x = 3, x = –1, x = 2 x = 3, x = 1, x = –2 x = –3, x = 1, x = –2

OpenStudy (anonymous):

just plug the numbers in and see if you get zero. For example, to test "1", you get 1 + 4 + 1 - 6 = 0, so x = 1 is a zero of the function. test the rest like this.

OpenStudy (anonymous):

Sum of the zeroes= -b/a=-4 Therefore, x= -3, x=1, x=-2 Or, -3+(-2)+1=-4 Therefore, option(d) is correct.

OpenStudy (anonymous):

Thanks a lot!!!